Kylan is four times as old as Ella. Quentin is one year younger than Kylan and five years older than Ella.

Answer by ikleyn(52787)

.
Kylan is four times as old as Ella. Quentin is one year younger than Kylan and five years older than Ella.
~~~~~~~~~~~~~~~~~

Let x be the Ella's age, in years.

Then Kylan age is 4x years.


Quentin age is, from one side, 4x-1 years,
and from the other side, x+5 years.

It gives this equation

    4x - 1 = x + 5.


Solve it, step by step

    4x - x = 5 + 1

      3x   =   6

       x   =   6/3 = 2.


ANSWER.  Ella is 2 years old;  Kylan is 4*2 = 8 years old;  Quentin is 8-1 = 7 years old.

Solved.